This question is "Least Common Multiple". I couldn't find a tab for it:
12=
35=
42=

Answer by jim_thompson5910(35256)

List the prime factorization of each number

12=2*2*3
35=5*7
42=2*3*7


Now circle the unique (and most frequent) factors. These factors are the two "2" terms in 12, any 3, a 5 and a 7. Multiply these all together to get 2*2*3*5*7=4*15*7 = 60*7 = 420


So the LCM of 12, 35, and 42 is 420


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim